# Automate Strategy Finding with LLM in Quant Investment

### v3

@@ Abstract
We present a novel three-stage framework leveraging Large Language Models (LLMs) within a risk-aware multi-agent system for automate strategy finding in quantitative finance. Our approach addresses the brittleness of traditional deep learning models in financial applications by: employing prompt-engineered LLMs to generate executable alpha factor candidates across diverse financial data, implementing multimodal agent-based evaluation that filters factors based on market status, predictive quality while maintaining category balance, and deploying dynamic weight optimization that adapts to market conditions. Experimental results demonstrate the robust performance of the strategy in Chinese & US market regimes compared to established benchmarks. Our work extends LLMs capabilities to quantitative trading, providing a scalable architecture for financial signal extraction and portfolio construction. The overall framework significantly outperforms all benchmarks with 53.17% cumulative return on SSE50  [^1] The SSE 50 Index tracks the performance of the 50 most influential large-cap blue-chip stocks on the Shanghai Stock Exchange. (Jan 2023 to Jan 2024), demonstrating superior risk-adjusted performance and downside protection on the market.
@@
Zhizhuo Kou1  Holam Yu2  Junyu Luo3  Jingshu Peng1  Xujia Li1 Chengzhong Liu1Juntao Dai3Lei Chen1,2Sirui Han1Yike Guo1 1The Hong Kong University of Science and Technology, Hong Kong SAR 2The Hong Kong University of Science and Technology (Guangzhou), Guangzhou, China 3Peking University, Beijing, China {alankou, leexujia, leichen, siruihan, yikeguo}@ust.hk hyu901@connect.hkust-gz.edu.cn, luojunyu@stu.pku.edu.cn jtd.acad@gmail.com, {jpengab,chengzhong.liu}@connect.ust.hk
@@ 1 Introduction
LLMs have evolved from supporting analytical tools to active participants in financial decision-making. For example, BloombergGPT demonstrates superior performance in parsing market sentiment and answering domain-specific questions Wu et al. (2023). Research shows that LLMs effectively generate trading actions by contextualizing price trends with news and earnings reports Ding et al. (2023). Concurrently, multi-agent systems offer powerful approaches to portfolio optimization through decentralized interaction. The Multi-Agent Portfolio System demonstrates the portfolio managed by agents have achieved well-diversified returns with improved risk-adjusted performance Lee et al. (2025). These frameworks capture complex market dynamics such as information sharing and strategic arbitrage that single-agent models cannot address Spooner and Savani (2020). The integration of these technologies creates sophisticated financial environments where LLM-enhanced agents demonstrate adaptive behavior. StockAgent exemplifies this approach with LLM-powered agents mimic diverse investor personas responding to market events Zhang et al. (2024a). Hierarchical structures such as FinCon organize agents in manager-analyst relationships, facilitating collaboration through natural language communication Yu et al. (2024). This convergence heralds a future of intelligent, distributed financial decision making that combines data-driven learning with human-like reasoning capabilities for more robust investment strategies Yu et al. (2023b).
@@
The proposed framework demonstrates versatility across various asset classes, enhancing its utility and practical effectiveness. To support future research and ensure reproducibility, we make source code publicly available at https://anonymous.4open.science/r/Automate-Strategy-Finding-with-LLM-in-Quant-investment-63C5.
@@ 3.3 Multimodal Multi-Agent Evaluation
Sharpe $\text{Sharpe Ratio}=\frac{\mathbb{E}[R(\alpha^{(t)})-R_{f}]}{\sqrt{\text{Var}[R(\alpha^{(t)})]}}$ (9)
@@ 3.4 Weight Optimization
This methodology establishes a robust framework for predicting future yields based on historical alpha values, demonstrating the efficacy of deep learning techniques in optimizing alpha weights and enhancing investment strategy performance. Algorithm 2 provides a formal specification of our complete framework, illustrating the logical flow from multimodal data processing through multi-agent evaluation to final weight optimization.
@@ 4.2 Multimodal Knowledge Extraction and Adaptive Alpha Discovery
We implement a prompt architecture (Figure 2) that incorporates multimodal market information into LLMs for comprehensive knowledge extraction and seed alpha selection. This architecture integrates textual financial sentiment data, numerical company financial statements, and visual trading charts to provide holistic market analysis. Our contextual analysis mechanism dynamically adjusts parameters based on prevailing market trends and sector performance. Experimental validation demonstrates the framework’s adaptability across varying market conditions. In Case 1, analyzing SSE50 data from Jan 2023 to Dec 2023, the model selected momentum and volume-based indicators such as price momentum, RSI, and MACD. Conversely, in Case 2, when processing the time window from Jan 2022 to Dec 2022, the model prioritized volatility and economic factors, including ATR, Bollinger Bands, and gross profit indicators.
@@ 4.3 Comparative Performance Against Traditional Alpha Factories
Momentum Mean Reversion Volatility Fundamental Growth
@@ 4.4 Performance Evaluation of the Integrated Alpha Framework in Benchmark Comparison
Table 3 presents an example combination of 12 alphas generated by our framework, evaluated on the SSE50 constituent stock set. The table details the seed alphas selected by the LLM from each category, along with their respective weights and IC 1 values. The weight combination IC 1 value is quite high as -0.0587. Although some of the seed alphas exhibit relatively low IC 1 values individually, their removal results in a significant drop in the retrained combination weight, indicating their critical role in the overall performance. For example, if we remove alpha #6 the weight combination will drop to -0.055; once we remove the alpha #11, the weight combination will drop to merely 0.0491. This suggests that the seed alpha set selected by the LLM synergizes effectively, providing robust predictive power Zhang et al. (2020). To address the question of whether our overall strategy, incorporating the LLM-driven framework and confidence scoring, can consistently beat the market, we conducted a backtest using a straightforward investment approach during the period from Jan 2023 to Jan 2024 (training: Jan 2021-Jun 2022; validation: Jul 2022-Dec 2022), on the SSE50 dataset we conduct the framework to construct the portfolio ( A.8). The net worth progression of the respective strategies over the testing period is shown in Figure 3, and the comparison matrix shows in Table 4. Despite not explicitly optimizing for absolute returns, our framework demonstrates impressive performance in the backtest, achieving the highest profit compared to other methods. Our cumulative return for 2023 backtest comes to $53.17\%$ positive and meanwhile the index performance is -11.73%. We also do the compression with 4 advanced statistical methods and 2 LLM based methods. The framework strategy substantially outperforms all competitors with the highest return (53.17%), best risk-adjusted metrics (Sharpe: 0.287, Sortino 222 Sortino ratio computing excess return per unit of downside deviation, evaluating only negative return volatility.: 0.208, Calmar  [^3] Calmar ratio divides annualized return by maximum drawdown, measuring reward relative to risk.: 1.052), and lowest volatility (0.762%). LLM based mathods FinCon and SEP show moderate success (22.47%, 17.89%), while conventional machine learning methodologies yield marginal returns (XGBoost 9.53%). Notably, all evaluated strategies surpass the SSE50 benchmark, which exhibits negative performance (-13.22%). This approach demonstrates the efficacy of our LLM-driven framework consistently outperforming market benchmarks through dynamic adaptation to changing conditions.
@@
Strategy Final Return (%) Sharpe Ratio Volatility (%) Sortino Ratio Calmar Ratio
@@ 4.5 Framework Robustness Across Time Periods and Markets
The empirical evidence substantiates both the temporal persistence and cross-market robustness of our findings, with notably beneficial countercyclical characteristics emerging during periods of market distress. The framework demonstrates consistent alpha generation across diverse market architectures, regulatory frameworks, and investor behavioral patterns, establishing its broad applicability. Specifically, the model maintains stable outperformance trajectories through bullish, bearish, and range-bound market conditions (detailed in Appendix A.9), thereby validating its structural integrity and adaptability to varying macroeconomic environments.
@@ 4.6 Ablation Study
Model Configuration IC (In-Sample) IC (Out-of-Sample) Sharpe Ratio
@@
Model Configuration Bull Market Bear Market Sideways Market
@@ 5 Related Work
The development of general-domain LLMs has catalyzed interest in Finance LLMs (Fin-LLMs), although this specialized domain remains nascent Novy-Marx (2015); Yang et al. (2023); Zhao et al. (2023). Open-source LLMs such as LLaMA Touvron et al. (2023), BLOOM W. and et al. (2023), and Flan-T5 W. and et al. (2022) provide flexibility but may underperform proprietary alternatives. Fine-tuned financial LLMs demonstrate enhanced domain-specific comprehension, yet their generative performance indicates the need for improved domain-specific datasets Lewis and et al. (2020); Koa et al. (2024).
Multimodal LLMs have shown significant potential in investment contexts by processing diverse data types L et al. (2023), developing strategies that mitigate market volatility K et al. (2024), and analyzing textual data to gauge investor sentiment Zhao et al. (2024). Multi-agent LLM systems enhance market analysis capabilities by leveraging vast datasets to interpret financial reports and market sentiment Zhang et al. (2024b), simulating various market scenarios Talebirad and Nadiri (2023), and facilitating parallel testing of diverse strategies Wang et al. (2024). Implementation raises important considerations regarding transparency, accountability, and bias mitigation Yu and et al. (2024); Mundhenk et al. (2021).
@@ 6 Conclusion
In this paper, we proposed a novel quantitative investment framework integrating LLMs and multi-Agent architectures to address instability in traditional approaches. Our system generates diversified alpha factors from multimodal financial data, constructs risk-calibrated trading agents, and employs a deep learning mechanism for dynamic agent weighting based on market conditions.
Experimental results confirm the framework’s effectiveness across Chinese and US markets, our framework demonstrates significant outperformance versus SOTA alpha generation methods and benchmark indices across key financial metrics. This work successfully extends LLM capabilities to quantitative trading, creating a scalable, adaptive architecture for financial signal extraction that functions effectively without human intervention across diverse market regimes.
@@ 7 Limitations
Our framework presents several significant limitations. First, system efficacy is contingent upon input document quality, potentially perpetuating inherent biases Deb et al. (2017); Ashok et al. (2018). Second, LLM-generated alphas occasionally lack the financial intuition characteristic of human analysts, resulting in theoretically sound but practically infeasible factors Tuarob et al. (2017). Third, our multi-agent evaluation methodology presupposes persistent historical relationships between market conditions and alpha performance—an assumption that may prove tenuous during market regime shifts. Finally, our validation efforts have primarily targeted equity markets, with cross-asset applicability requiring additional empirical investigation. Future research directions should address these constraints through exploring Mixture of Experts (MoE) architectures to improve learning efficiency Masoudnia and Ebrahimpour (2014), adaptive agent architectures, transfer learning methodologies Wang et al. (2023), and computationally efficient implementations.
@@ A.1 Sample of Seed Alpha
Document name Author
Financial Time Series Prediction Using Mixture of Experts M. Serdar Yumlu, Fikret S. Gurgen, and Nesrin Okay
Model-Free Implied Dependence and the Cross-Section of Returns Koen Inghelbrecht, Gertjan Verdickt, Daniël Linders, and Yong Xie
Dynamic Graph-based Deep Reinforcement Learning with Long and Short-term Relation Modeling for Portfolio Optimization Pengqian Yu, Joon Sern Lee, Ilya Kulyatin, Zekun Shi, Sakyasingha Dasgupta
Can Large Language Models Beat Wall Street? Unveiling the Potential of AI in Stock Selection Georgios Fatouros, Konstantinos Metaxas, John Soldatos, Dimosthenis Kyriazis
Mastering Stock Markets with Efficient Mixture of Diversified Trading Experts Shuo Sun, Xinrun Wang, Wanqi Xue, Xiaoxuan Lou, and Bo An
101 Formulaic Alphas Zura Kakushadze
A Multimodal Foundation Agent for Financial Trading: Tool-Augmented, Diversified, and Generalist Wentao Zhang, Lingxuan Zhao, Haochong Xia, Shuo Sun, Jiaze Sun, Molei Qin, Xinyi Li, Yuqing Zhao, Yilei Zhao, Xinyu Cai, Longtao Zheng, Xinrun Wang, Bo An
Causal Factor Investing: Can Factor Investing Become Scientific? Marcos Lopez de Prado
Alpha-GPT: Human-AI Interactive Alpha Mining for Quantitative Investment Saizhuo Wang, Hang Yuan, Leon Zhou, Lionel M. Ni, Heung-Yeung Shum, Jian Guo
AlphaPortfolio: Direct Construction Through Deep Reinforcement Learning and Interpretable AI Lin William Cong, Ke Tang, Jingyuan Wang, Yang Zhang
FinGPT: Open-Source Financial Large Language Models Hongyang (Bruce) Yang, Xiao-Yang Liu, Christina Dan Wang
@@
Chande Momentum Oscillator (CMO) (SUM(IF(CLOSE - DELAY(CLOSE, 1) > 0, CLOSE - DELAY(CLOSE, 1), 0), 14) - SUM(IF(CLOSE - DELAY(CLOSE, 1) $<$ 0, DELAY(CLOSE, 1) - CLOSE, 0), 14)) / (SUM(IF(CLOSE - DELAY(CLOSE, 1) > 0, CLOSE - DELAY(CLOSE, 1), 0), 14) + SUM(IF(CLOSE - DELAY(CLOSE, 1) $<$ 0, DELAY(CLOSE, 1) - CLOSE, 0), 14)) * 100

### v4

@@ Abstract
We present a novel three-stage framework leveraging Large Language Models (LLMs) within a risk-aware multi-agent system for automate strategy finding in quantitative finance. Our approach addresses the brittleness of traditional deep learning models in financial applications by: employing prompt-engineered LLMs to generate executable alpha factor candidates across diverse financial data, implementing multimodal agent-based evaluation that filters factors based on market status, predictive quality while maintaining category balance, and deploying dynamic weight optimization that adapts to market conditions. Experimental results demonstrate the robust performance of the strategy in Chinese & US market regimes compared to established benchmarks. Our work extends LLMs capabilities to quantitative trading, providing a scalable architecture for financial signal extraction and portfolio construction. The overall framework significantly outperforms all benchmarks with 53.17% cumulative return on SSE50  [^1] The SSE 50 Index tracks the performance of the 50 most influential large-cap blue-chip stocks on the Shanghai Stock Exchange. (Jan 2023 to Jan 2024), demonstrating superior risk-adjusted performance and downside protection on the market. Our protocol code are available at https://github.com/kouzhizhuo/Automate-Strategy-Finding-with-LLM-in-Quant-investment.
@@
Zhizhuo Kou1, Holam Yu2, Junyu Luo3, Jingshu Peng1, Xujia Li1, Chengzhong Liu1, Juntao Dai3, Lei Chen1, 2, Sirui Han1thanks: Corresponding author. thanks: Project leader Yike Guo [^1] footnotemark: 1, 1 The Hong Kong University of Science and Technology 2 The Hong Kong University of Science and Technology (Guangzhou) 3 Peking University zkouaa@connect.ust.hk {siruihan, yikeguo}@ust.hk
@@ 1 Introduction
LLMs have evolved from supporting analytical tools to active participants in financial decision-making Luo et al. (2025c). For example, BloombergGPT demonstrates superior performance in parsing market sentiment and answering domain-specific questions Wu et al. (2023). Research shows that LLMs effectively generate trading actions by contextualizing price trends with news and earnings reports Ding et al. (2023). Concurrently, multi-agent systems offer powerful approaches to portfolio optimization through decentralized interaction. The Multi-Agent Portfolio System demonstrates the portfolio managed by agents have achieved well-diversified returns with improved risk-adjusted performance Lee et al. (2025). These frameworks capture complex market dynamics such as information sharing and strategic arbitrage that single-agent models cannot address Spooner and Savani (2020). The integration of these technologies creates sophisticated financial environments where LLM-enhanced agents demonstrate adaptive behavior. StockAgent exemplifies this approach with LLM-powered agents mimic diverse investor personas responding to market events Zhang et al. (2024a). Hierarchical structures such as FinCon organize agents in manager-analyst relationships, facilitating collaboration through natural language communication Yu et al. (2024). This convergence heralds a future of intelligent, distributed financial decision making that combines data-driven learning with human-like reasoning capabilities for more robust investment strategies Yu et al. (2023b).
@@
The proposed framework demonstrates versatility across various asset classes, enhancing its utility and practical effectiveness. To support future research and ensure reproducibility, we make source code publicly available at https://github.com/kouzhizhuo/Automate-Strategy-Finding-with-LLM-in-Quant-investment.
@@ 3.3 Multimodal Multi-Agent Evaluation
$\text{Sharpe Ratio}=\frac{\mathbb{E}[R(\alpha^{(t)})-R_{f}]}{\sqrt{\text{Var}[R(\alpha^{(t)})]}}$ (9)
@@ 3.4 Weight Optimization
This methodology establishes a robust framework for predicting future yields based on historical alpha values, demonstrating the efficacy of deep learning techniques in optimizing alpha weights and enhancing investment strategy performance. Algorithm LABEL:overallago provides a formal specification of our complete framework, illustrating the logical flow from multimodal data processing through multi-agent evaluation to final weight optimization.
@@ 4.2 Multimodal Knowledge Extraction and Adaptive Alpha Discovery
We implement a prompt architecture (Figure 2) that incorporates multimodal market information into LLMs for comprehensive knowledge extraction and seed alpha selection. This architecture integrates textual financial sentiment data, numerical company financial statements, and visual trading charts to provide holistic market analysis Luo et al. (2025b). Our contextual analysis mechanism dynamically adjusts parameters based on prevailing market trends and sector performance. Experimental validation demonstrates the framework’s adaptability across varying market conditions. In Case 1, analyzing SSE50 data from Jan 2023 to Dec 2023, the model selected momentum and volume-based indicators such as price momentum, RSI, and MACD. Conversely, in Case 2, when processing the time window from Jan 2022 to Dec 2022, the model prioritized volatility and economic factors, including ATR, Bollinger Bands, and gross profit indicators.
@@ 4.3 Comparative Performance Against Traditional Alpha Factories
Momentum
Mean
Reversion
Volatility Fundamental Growth
@@ 4.4 Performance Evaluation of the Integrated Alpha Framework in Benchmark Comparison
Table 3 presents an example combination of 12 alphas generated by our framework, evaluated on the SSE50 constituent stock set. The table details the seed alphas selected by the LLM from each category, along with their respective weights and IC 1 values. The weight combination IC 1 value is quite high as -0.0587. Although some of the seed alphas exhibit relatively low IC 1 values individually, their removal results in a significant drop in the retrained combination weight, indicating their critical role in the overall performance. For example, if we remove alpha #6 the weight combination will drop to -0.055; once we remove the alpha #11, the weight combination will drop to merely 0.0491. This suggests that the seed alpha set selected by the LLM synergizes effectively, providing robust predictive power Zhang et al. (2020). To address the question of whether our overall strategy, incorporating the LLM-driven framework and confidence scoring, can consistently beat the market, we conducted a backtest using a straightforward investment approach during the period from Jan 2023 to Jan 2024 (training: Jan 2021-Jun 2022; validation: Jul 2022-Dec 2022), on the SSE50 dataset we conduct the framework to construct the portfolio ( A.8). The net worth progression of the respective strategies over the testing period is shown in Figure 3, and the comparison matrix shows in Table 4. Despite not explicitly optimizing for absolute returns, our framework demonstrates impressive performance in the backtest, achieving the highest profit compared to other methods. Our cumulative return for 2023 backtest comes to $53.17\%$ positive and meanwhile the index performance is -11.73%. We also do the compression with 4 advanced statistical methods and 2 LLM based methods. The framework strategy substantially outperforms all competitors with the highest return (53.17%), best risk-adjusted metrics (Sharpe: 0.287, Sortino  [^2] Sortino ratio computing excess return per unit of downside deviation, evaluating only negative return volatility.: 0.208, Calmar  [^3] Calmar ratio divides annualized return by maximum drawdown, measuring reward relative to risk.: 1.052), and lowest volatility (0.762%). LLM based mathods FinCon and SEP show moderate success (22.47%, 17.89%), while conventional machine learning methodologies yield marginal returns (XGBoost 9.53%). Notably, all evaluated strategies surpass the SSE50 benchmark, which exhibits negative performance (-13.22%). This approach demonstrates the efficacy of our LLM-driven framework consistently outperforming market benchmarks through dynamic adaptation to changing conditions.
@@
Strategy
Final
Return (%)
Sharpe
Ratio
Volatility
Sortino
Ratio
Calmar
Ratio
@@ 4.5 Framework Robustness Across Time Periods and Markets
The empirical evidence substantiates both the temporal persistence and cross-market robustness of our findings, with notably beneficial countercyclical characteristics emerging during periods of market distress. The framework demonstrates consistent alpha generation across diverse market architectures, regulatory frameworks, and investor behavioral patterns, establishing its broad applicability. Specifically, the model maintains stable outperformance trajectories through bullish, bearish, and range-bound market conditions, thereby validating its structural integrity and adaptability to varying macroeconomic environments.
@@ 4.6 Ablation Study
Model Configuration
IC
(In-Sample)
IC
(Out-of-Sample)
Sharpe
Ratio
@@
Model Configuration
Bull
Market
Bear
Market
Sideways
Market
@@ 5 Related Work
The development of general-domain LLMs has catalyzed interest in Finance LLMs (Fin-LLMs), although this specialized domain remains nascent Novy-Marx (2015); Yang et al. (2023); Zhao et al. (2023). Open-source LLMs such as LLaMA Touvron et al. (2023), BLOOM W. and et al. (2023), and Flan-T5 W. and et al. (2022) provide flexibility but may underperform proprietary alternatives. Fine-tuned financial LLMs demonstrate enhanced domain-specific comprehension, yet their generative performance indicates the need for improved domain-specific datasets Lewis and et al. (2020); Koa et al. (2024); Wen et al. (2025).
Multimodal LLMs have shown significant potential in investment contexts by processing diverse data types L et al. (2023); Luo et al. (2025a), developing strategies that mitigate market volatility K et al. (2024), and analyzing textual data to gauge investor sentiment Zhao et al. (2024). Multi-agent LLM systems enhance market analysis capabilities by leveraging vast datasets to interpret financial reports and market sentiment Zhang et al. (2024b), simulating various market scenarios Talebirad and Nadiri (2023), and facilitating parallel testing of diverse strategies Wang et al. (2024). Implementation raises important considerations regarding transparency, accountability, and bias mitigation Yu and et al. (2024); Mundhenk et al. (2021).
@@ 6 Conclusion
In this paper we proposed a novel quantitative investment framework integrating LLMs and multi-Agent architectures to address instability in traditional approaches. Our system generates diversified alpha factors from multimodal financial data, constructs risk-calibrated trading agents, and employs a deep learning mechanism for dynamic agent weighting based on market conditions.
Experimental results confirm the framework’s effectiveness across CN and US markets, our framework demonstrates significant outperformance versus SOTA alpha generation methods and benchmark indices across key financial metrics. This work successfully extends LLM capabilities to quantitative trading, creating a scalable, adaptive architecture for financial signal extraction that functions effectively without human intervention.
@@ 7 Limitations
Our framework presents several significant limitations. First, system efficacy is contingent upon input document quality, potentially perpetuating inherent biases Deb et al. (2017); Ashok et al. (2018). Second, LLM-generated alphas occasionally lack the financial intuition characteristic of human analysts, resulting in theoretically sound but practically infeasible factors Tuarob et al. (2017). Third, our multi-agent evaluation methodology presupposes persistent historical relationships between market conditions and alpha performance, an assumption that may prove tenuous during market regime shifts. Finally, our validation efforts have primarily targeted equity markets, with cross-asset applicability requiring additional empirical investigation. Future research directions should address these constraints through exploring Mixture of Experts (MoE) architectures to improve learning efficiency Masoudnia and Ebrahimpour (2014), adaptive agent architectures, transfer learning methodologies Wang et al. (2023), related regulation requirements Han and Xi (2020), and computationally efficient implementations.
8 Acknowledgments
This paper is partially supported by grants from the HKUST Start-up Fund (R9911), Theme-based Research Scheme grant (T45-205/21-N), the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government, and the research funding under HKUST-DXM AI for Finance Joint Laboratory (DXM25EG01). Also supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China (NSFC) under Grant No. U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No. 2024A0505040012, the Hong Kong RGC GRF Project 16213620, RIF Project R6020-19, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Key Areas Special Project of Guangdong Provincial Universities 2024ZDZX1006, Guangdong Province Science and Technology Plan Project 2023A0505030011, Guangzhou municipality big data intelligence key lab, 2023A03J0012, Hong Kong ITC ITF grants MHX/078/21 and PRP/004/22FX, Zhujiang scholar program 2021JC02X170, Microsoft Research Asia Collaborative Research Grant, HKUST-Webank joint research lab and 2023 HKUST Shenzhen-Hong Kong Collaborative Innovation Institute Green Sustainability Special Fund, from Shui On Xintiandi and the InnoSpace GBA. The authors are grateful to anonymous reviewers for their efforts and insightful suggestions to improve this article.
@@ A.1 Sample of Seed Alpha
Chande Momentum Oscillator (CMO) (SUM(IF(CLOSE - DELAY(CLOSE, 1) > 0, CLOSE - DELAY(CLOSE, 1), 0), 14) - SUM(IF(CLOSE - DELAY(CLOSE, 1) << 0, DELAY(CLOSE, 1) - CLOSE, 0), 14)) / (SUM(IF(CLOSE - DELAY(CLOSE, 1) > 0, CLOSE - DELAY(CLOSE, 1), 0), 14) + SUM(IF(CLOSE - DELAY(CLOSE, 1) << 0, DELAY(CLOSE, 1) - CLOSE, 0), 14)) * 100